Evaluate the limit "\displaystyle{\lim\limits_{\theta \to 0} \dfrac{\sin{(\theta^2)}}{\theta}}" using the l'Hopital's Rule.


1
Expert's answer
2021-09-19T18:15:56-0400

Since we have an indeterminate form of type "\frac{0}{0}", we can apply the l'Hopital's rule.

"\lim_{\theta \to 0} \frac{\sin{\theta^2}}{\theta}=\lim_{\theta \to 0} \frac{\frac{d}{d\theta}(\sin{\theta^2})}{\frac{d}{d\theta}(\theta)}=\lim_{\theta \to 0}2 \theta cos{\theta^2}=0\\\\
\text{Hence }\\
\lim_{\theta \to 0} \frac{\sin{\theta^2}}{\theta}=0"